Write x˄1/6*y˄1/7*z˄1/4 using rational exponents as a single expression
Answer by lwsshak3(11628)

Write x˄1/6*y˄1/7*z˄1/4 using rational exponents as a single expression.
...
f(x)=x˄1/6*y˄1/7*z˄1/4
LCD for fractional exponents=84
1/6=14/84
1/7=12/84
1/4=21/84
..
f(x)=x˄1/6*y˄1/7*z˄1/4
f(x)=x^14/84*y^12/84*z^21/84
f(x)=(xyz)^47/84